A. Direction: Find the sample proportion. (Show your complete solution.)

1. n = 550, x = 308

2. n = 480, x = 168

3. n = 620, x = 248

4. n = 740, x = 259

5. n = 1,120, x = 616


1
Expert's answer
2021-05-21T07:09:26-0400

"P=\frac{x}{n}"

1. "P=\frac{308}{550}=0.56"

2. "P=\frac{168}{480}=0.35"

3. "P=\frac{248}{620}=0.4"

4. "P=\frac{259}{740}=0.35"

5. "P=\frac{616}{1120}=0.55"